# AUTHOR ONE et al

### v1

@@ 3 Considerations for selecting a randomization ratio
As shown in Section 2, power of the logrank test is maximized in event-driven survival trials under PH when the number of events is balanced across treatment arms. Since pivotal trials typically target a certain level of statistical power (e.g., 80% or 90%), a randomization ratio (RR) that is sub-optimal with respect to power may require more events to be observed, which may in turn impact trial duration or sample size. We now describe in more detail these design aspects along with other considerations when selecting a RR.
@@ 3.1 Accrual and trial duration
Investigators often cite accelerating patient recruitment as one of the reasons for implementing unequal randomization.5, 6 The assumption is that patients are more likely to participate in a randomized trial when their odds of receiving the experimental therapy are greater. This may be especially true if the control therapy is unattractive (e.g., placebo).
@@
In addition, it is unclear whether randomizing more patients to the experimental arm actually accelerates recruitment.5 The reality is that RR is just one of many factors that patients consider when deciding whether or not to participate in a trial. Other important factors include duration of treatment, mode of administration, out of pocket costs, travel time, and home visits.16 From a systems perspective, insufficient infrastructure and restrictive eligibility criteria also present significant barriers to clinical trial participation.17, 18
@@ 3.2 Sample size
Sponsors of survival trials have flexibility when it comes to deciding the overall sample size of a trial. According to Schoenfeld’s approximation, a trial under PH can achieve the target level of power under various number of patients, so long as the target number of events is observed. While sponsors could theoretically calculate number of patients with the intention of achieving balance of events at the time of analysis (thereby making the most efficient use of events), the standard practice is to target an event-patient ratio ($d/n$) that balances resource constraints such as the available pool of patients, trial costs, and trial timeline.
Therefore, a more realistic consideration is as follows: If employing unequal randomization would increase trial duration (compared to 1:1 randomization and assuming fixed accrual rate and sample size), how many additional patients would need to be recruited to avoid any delay in trial readout? Doing so would impact recruitment and trial costs, but it would also even the playing field with respect to trial duration. We shall explore “edge cases” such as this in Section 4.
@@ 3.3 Balance in baseline characteristics
A related consideration to sample size is the balance in baseline characteristics. Unequal randomization is more likely to result in an imbalance in baseline characteristics between treatment arms and thus chance bias.19 This concern applies more to small- to medium-sized trials as well as subgroups within a trial, since they carry a higher risk of imbalance to begin with.
To illustrate the impact of unequal randomization, let $X$ be a baseline covariate with standard deviation $\sigma$. A trial with $n_{\pi}$ patients that randomizes $\pi\times 100\%$ patients to the experimental arm and $(1-\pi)\times 100\%$ patients to the control arm will have the following variance for the difference in average baseline:
$\mbox{Var}(\bar{X}_{1}-\bar{X}_{0})=\frac{\sigma}{n_{\pi}}\cdot\frac{1}{\pi(1-\pi)}.$
This quantity is directly associated to the risk of baseline imbalance. Furthermore, it is equivalent to the variance for the difference in average baseline in a trial employing equal randomization, but with sample size
@@
For example, a similar amount of baseline imbalance can be expected in a 2:1 randomized trial with 180 patients as in a 1:1 randomized trial with 160 patients.
@@ 3.4 Secondary estimands
The ICH E9(R1) Addendum introduced the estimands framework as a structured approach to formulating clinical trial objectives.20 Under this framework, an estimand (i.e., a precise description of the treatment effect reflecting the clinical question posed by a clinical trial objective) consists of five attributes: population, treatment, endpoint, intercurrent events and strategy, and population-level summary. Below are a few examples of estimand attributes that may define important secondary objectives in oncology trials. These examples should not be considered as collectively defining a single estimand, but rather as individual, defining attributes that differentiate a secondary objective from the primary objective.
@@
For any given secondary objective and chosen estimator, what constitutes as sufficient or mature follow-up needs to be defined.21, 22, 23, 24 For example, one might consider a Kaplan-Meier estimate of the 2-year survival probability to be mature provided that the width of its confidence interval is within a certain range (“precision” and “reliability”), or that the estimate will change little given longer follow-up (“stability”). In other cases, defining mature follow-up may be difficult (e.g., overall survival in indolent cancers25).
@@ 3.5 Other
If a treatment arm is substantially more expensive than the other, then it may be desirable to assign more patients to the less expensive intervention treatment.26, 27 For example, CAR T-cell therapies carry high costs because of their collection and manufacturing process. Meanwhile, therapies that were recently approved and are currently under exclusivity may be expensive for sponsors of a new trial to obtain as the control therapy.
@@
Randomizing more patients to experimental therapy may benefit enrolled patients by prolonging their time-to-event (improved individual ethics). However, it may also prolong trial duration, in which case patients not enrolled in the trial would have to wait longer for drug approval and access (worsened collective ethics). Other factors that may tip the argument in favor of individual or collective ethics include disease prevalence, disease severity, strength of early phase results, and safety profile of the experimental therapy.28
@@ 4 Case studies
We now consider various trial designs employing 1:1 randomization and alternative designs that differ in randomization ratio, accrual duration, trial duration, and/or sample size. As per usual in trial design, power is fixed at some desirable level, i.e., 80%.
@@ 4.1 Checkmate-017
Checkmate-017 was a randomized, open-label phase 3 trial that compared nivolumab vs. docetaxel in patients with advanced squamous-cell non–small-cell lung cancer.29 The trial planned to enroll 264 patients over a period of 12 months, to randomize the patients 1:1 to nivolumab and docetaxel, and to follow for overall survival until 189 deaths were observed. Given two-sided alpha 0.04 (0.01 was reserved for the second primary endpoint overall response rate) and assuming exponential survival distributions with medians 11.4 and 7.0 months, the trial was projected to take 24 months to complete and to have 90% statistical power to detect a target hazard ratio of $7.0/11.4=0.61$.
For the purpose of this case-study, we assume that Checkmate-017’s protocol follows similar assumptions as above (1:1 randomization; exponential survival distributions with medians 11.4 and 7.0 months; 22 patients/month enrollment rate; 0.72 event-patient ratio), but with the following differences and additional assumptions: trial is powered at 80% given two-sided alpha 0.05, enrollment follows a uniform distribution, and dropout follows an exponential distribution so that patients have a 5% probability of dropping out every 12 months. Consequently, this trial design plans to enroll 186 patients over a period of $\sim$8.5 months and to follow the patients until 133 deaths are observed for a total trial duration of $\sim$21.7 months (CM-017 in Table 1).
@@
  • Prolonged trial: assume unequal randomization will have no impact on enrollment rate (22 patients/month). Under the alternative hypothesis, patients receiving experimental therapy will have improved survival. Thus, randomizing patients 3:2 or 2:1 may result in more time needed to observe the required number of events.
  • Accelerated accrual: assume unequal randomization will accelerate enrollment to such a rate that the original trial timeline of $\sim$21.7 months is unaffected.
  • Increased enrollment: assume unequal randomization will have no impact on enrollment rate (22 patients/month), but additional patients are enrolled at such numbers that the original trial timeline of $\sim$21.7 months is unaffected.
@@
Note that event sizes in Table 1 were determined empirically via simulations. Had Rubinstein’s equation been used instead, similar event sizes would have been obtained (134 for 3:2, 141 for 2:1). However, had Schoenfeld’s equation been used, notably larger event sizes would have been suggested to maintain 80% power (138 for 3:2, 149 for 2:1), which in turn would have exaggerated operating characteristics of the alternative trial designs (i.e., longer trial duration, larger sample size or higher enrollment rates needed to avoid delay). These results support our recommendation at the end of Section 2.4 to avoid using Schoenfeld’s approximation and to instead use Rubinstein’s approximation or empirical simulations when evaluating trial designs under unequal randomization.
@@ 4.2 More general cases
We now extend our exercise to cover a wider range of scenarios: hazard ratio (HR) 0.5, 0.6, 0.7, and 0.8; control median (CM) survival 6, 12, and 24 months; and event-patient ratio ($d/n$) 0.5, 0.6, 0.7, and 0.8. For each “original” trial design employing 1:1 randomization and defined by a unique triplet (HR, CM, $d/n$), we use the same combination of event size ($d$), sample size ($n$), and accrual rate as was done for the simulations described in Section 2.4. We assume that dropout follows an exponential distribution such that patients have a 1% probability of dropping out every 12 months. Finally, we identify six alternative trial designs by following the approach in Section 4 (“3:2, Prolonged trial”, “3:2, Accelerated accrual”, etc.). Results across all scenarios are summarized in Table 2. (For case-by-case results, see Tables S1–S3.)
Table 2: Comparing number of events and trial duration under 3:2 and 2:1 versus 1:1 randomization in trials with 80% power and various hazard ratios (HR), control medians (CM), and event-patient ratios ($d/n$). Since under the alternative hypothesis, unequal randomization (without increasing accrual rate or number of patients) leads to longer trials, we also consider the required change in accrual rate (without increasing number of patients) or number of patients (without increasing accrual rate) to avoid any trial delay.
@@
No. of events Minor difference (<$10\%$) Minor difference (<$10\%$), except when treatment effect is weaker and event-patient ratio is large (HR $\geq 0.7$, $d/n\geq 0.8$). In this case, no. of events increases >10%.
@@
Trial duration assuming same accrual rate and no. of patients Minor increase (<$10\%$), except when treatment effect is strong, survival is long, and event-patient ratio is large (HR $\leq 0.5$, CM $\geq 12$, $d/n\geq 0.8$). In this case, trial duration increases >10%. Minor increase (<$10\%$) when event-patient ratio is small ($d/n\leq 0.6$). Increase >$10\%$ otherwise.
@@
— Accrual rate to avoid increase in trial duration Moderate increase (<$25\%$) when survival is short or event-patient ratio is small (CM $\leq 6$ or $d/n\leq 0.6$). Increase >$25\%$ otherwise. Increase >$25\%$, except when survival is short and event-patient ratio is small (CM $\leq 6$, $d/n\leq 0.6$). In this case, accrual rate increase is moderate (<25%).
@@
— No. of patients to avoid increase in trial duration Minor increase (<$10\%$) Increase >$10\%$, except when treatment effect is strong and event-patient ratio is small (HR $\leq 0.6$, $d/n\leq 0.7$). In this case, increase no. of patients is minor (<10%).
@@
Compared to 1:1 randomization, 3:2 randomization requires similar number of events to achieve 80% power. In most cases, 3:2 randomization leads to a minor increase in trial duration (assuming the same accrual rate and number of patients), which can be avoided by a moderate increase in accrual rate or minor increase in number of patients. A notable exception is when survival is long and event-patient ratio is large (CM $\geq 12$, $d/n\geq 0.7$). In this case, a >25% increase in accrual rate is needed to avoid any trial delay. Trial delay may even be impossible to avoid in some cases; when event-patient ratio is 0.8, no amount of increase in accrual rate may offset the slower rate of accumulating events under 3:2 randomization.
@@
2:1 randomization also requires similar number of events as 1:1 randomization in many cases. However, there are some cases when 2:1 randomization requires a >10% increase in number of events (e.g., HR=0.8). 2:1 randomization leads to >10% increase in trial duration when event-patient ratio is large ($d/n\geq 0.7$). Trial delay can be difficult to avoid; in many cases, a >25% increase in accrual rate or >10% increase in number of patients is needed. And more often than 3:2 randomization, trial delay may be unavoidable. Thus, 2:1 randomization is more likely than 3:2 randomization to result in a notable increase in trial duration, and a greater increase in accrual rate or number of patients is needed to avoid it.
@@
It is worth noting that, in some cases where hazard ratio and event-patient ratio are small, we do see unequal randomization actually requiring less events than 1:1 randomization. However, the decrease in number of events is always small—2 at most. Therefore, consistent with our observation from Figure 3, 1:1 randomization may not maximize power, but it experiences little power loss compared to 3:2 and 2:1 randomization.
@@
The fact that Schoenfeld’s approximation is less accurate under unequal randomization may cause us to wonder whether we should stop conducting event-driven survival trials all together; part of the lure of this practice has been the thought that power depends on the underlying HR, but not some of the other design parameters. We see from Tables S1–S3 that indeed, for trial designs under unequal randomization, the required number of events (and therefore power) depends on the control median survival and event-patient ratio. However, we also see that this dependency is much less compared to the dependency on the hazard ratio (i.e., changes to $d$ as a function of CM or $d/n$ are much smaller compared to changes to $d$ as a function of HR).
@@
Our recommendation is therefore two-fold. First, randomized survival trials with logrank test as the primary analysis should continue to be event-driven. Second, for trials employing unequal randomization, the required number of events should be calculated either via Rubinstein’s approximation or empirical simulations, taking into careful consideration other design assumptions such as survival distribution in the control arm and event-patient ratio. Our recommendation for unequal randomization is consistent with the current practice for trials that anticipate non-proportional hazards (NPH) at the time of trial design; many of these trials remain event-driven, but consider potential NPH when calculating the required number of events or power (see Checkmate-27430 and Keynote-96631 for example).

### v2

@@ 3 Considerations for selecting a randomization ratio
As shown in Section 2, power of the LRT is maximized in event-driven survival trials under PH when the number of events is balanced across treatment arms. Since pivotal trials typically target a certain level of statistical power (e.g., 80% or 90%), a RR that is sub-optimal with respect to power may require more events to be observed, which may in turn impact trial duration or sample size. We now describe in more detail these design aspects along with other considerations when selecting a RR.
@@ 3.1 Accrual and trial duration
Investigators often cite accelerating patient recruitment as one of the reasons for implementing unequal randomization.5, 6 The assumption is that patients are more likely to participate in a randomized trial when their odds of receiving the experimental therapy are greater. This may be especially true if the control therapy is unattractive, such as a placebo.
@@
In addition, it is unclear whether randomizing more patients to the experimental arm actually accelerates recruitment.5 The reality is that RR is just one of many factors that patients consider when deciding whether or not to participate in a trial. Other important factors include duration of treatment, mode of administration, out-of-pocket costs, travel time, and home visits.18 From a systems perspective, insufficient infrastructure and restrictive eligibility criteria also present significant barriers to clinical trial participation.19, 20
@@ 3.2 Sample size
Sponsors of survival trials have flexibility when it comes to deciding the overall sample size of a trial. According to Schoenfeld’s approximation, a trial under PH can achieve the target level of power under various number of patients, so long as the target number of events is observed. While sponsors can theoretically calculate number of patients with the intention of achieving balance of events at the time of analysis (thereby making the most efficient use of events), the standard practice is to target an EPR that balances resource constraints such as the available pool of patients, trial costs, and trial timeline.
Therefore, a more realistic consideration is as follows: If employing unequal randomization would increase trial duration (compared to equal randomization and assuming fixed accrual rate and sample size), how many additional patients would need to be recruited to avoid any delay in trial readout? Doing so would impact recruitment and trial costs, but it would also even the playing field with respect to trial duration. We shall explore “edge cases” such as this in Section 4.
@@ 3.3 Balance in baseline characteristics
A related consideration to sample size is the balance in baseline characteristics. Unequal randomization is more likely to result in an imbalance in baseline characteristics between treatment arms and thus chance bias.21 This concern applies more to small- to medium-sized trials as well as subgroups within a trial, since they carry a higher risk of imbalance to begin with. Of course, if there are known factors that influence prognosis, then imbalance with respect to these prognostic factors can be prevented via stratified randomization.
To illustrate the potential impact of unequal randomization, let $X$ be a baseline covariate with standard deviation $\sigma$. A trial with $n_{\pi}$ patients that randomizes $\pi\times 100\%$ patients to the experimental arm and $(1-\pi)\times 100\%$ patients to the control arm will have the following variance for the difference in average baseline:
$\mbox{Var}(\bar{X}_{1}-\bar{X}_{0})=\frac{\sigma^{2}}{n_{\pi}}\cdot\frac{1}{\pi(1-\pi)}.$
This quantity is directly associated to the risk of baseline imbalance, since the larger the variance, the more likely it is for the trial to observe a large difference in average baseline.
This quantity is also equivalent to the variance for the difference in average baseline in a trial employing 1:1 randomization, but with sample size
@@
Thus, for example, a similar amount of baseline imbalance can be expected in a 2:1 randomized trial with 180 patients as in a 1:1 randomized trial with 160 patients.
@@ 3.4 Secondary estimands
The ICH E9(R1) Addendum introduced the estimands framework as a structured approach to formulating clinical trial objectives.22 Under this framework, an estimand (i.e., a precise description of the treatment effect reflecting the clinical question posed by a clinical trial objective) consists of five attributes: population, treatment, endpoint, intercurrent events and strategy, and population-level summary. Below are a few examples of estimand attributes that may define important secondary objectives in oncology trials. These examples should not be considered as collectively defining a single estimand, but rather as individual, defining attributes that differentiate a secondary objective from the primary objective.
@@
For any given secondary objective and chosen estimator, what constitutes as sufficient or mature follow-up needs to be defined.23, 24, 25, 26 For example, one might consider a Kaplan-Meier estimate of the 2-year survival probability to be mature provided that the width of its confidence interval is within a certain range (“precision” and “reliability”), or that the estimate will change little given longer follow-up (“stability”). In other cases, defining mature follow-up may be difficult (e.g., overall survival in indolent cancers27).
@@ 3.5 Other
If a treatment arm is substantially more expensive than the other, then it may be desirable to assign more patients to the less expensive intervention treatment.28, 29 For example, CAR T-cell therapies carry high costs because of their collection and manufacturing process. Meanwhile, therapies that were recently approved and are currently under exclusivity may be expensive for sponsors of a new trial to obtain as the control therapy.
@@
Randomizing more patients to experimental therapy may benefit enrolled patients by prolonging their time-to-event (improved individual ethics). However, it may also prolong trial duration, in which case patients not enrolled in the trial would have to wait longer for drug approval and access (worsened collective ethics). Other factors that may tip the argument in favor of individual or collective ethics include disease prevalence, disease severity, strength of early phase results, and safety profile of the experimental therapy.30
@@ 4 Case studies
We now consider various trial designs employing 1:1 randomization and alternative designs that differ in randomization ratio, accrual duration, trial duration, and/or sample size. Per usual in trial design, power is fixed at some level, in this case 80%.
@@ 4.1 Checkmate-017
Checkmate-017 was a randomized, open-label phase 3 trial that compared nivolumab vs. docetaxel in patients with advanced squamous-cell non–small-cell lung cancer.31 The trial planned to enroll 264 patients over a period of 12 months, to randomize the patients 1:1 to nivolumab and docetaxel, and to follow for overall survival until 189 deaths were observed. Given two-sided alpha 0.04 (0.01 was reserved for the second primary endpoint overall response rate) and assuming exponential survival distributions with medians 11.4 and 7.0 months, the trial was projected to take 24 months to complete and to have 90% statistical power to detect a target hazard ratio of $7.0/11.4=0.61$.
For the purpose of this case-study, we assume that Checkmate-017’s protocol follows similar assumptions as above (1:1 randomization, exponential survival distributions with medians 11.4 and 7.0 months, 22 patients/month enrollment rate, 0.72 event-patient ratio), but with the following differences and additional assumptions: trial is powered at 80% given two-sided alpha 0.05, enrollment follows a uniform distribution, and dropout follows an exponential distribution so that patients have a 5% probability of dropping out every 12 months. Consequently, this trial design plans to enroll 186 patients over a period of $\sim$8.5 months and to follow the patients until 133 deaths are observed for a total trial duration of $\sim$21.7 months (CM-017 in Table 1).
@@
  • Prolonged trial: assume unequal randomization will have no impact on enrollment rate (22 patients/month). Under the alternative hypothesis, patients receiving experimental therapy will have improved survival. Thus, randomizing patients 3:2 or 2:1 may result in more time needed to observe the required number of events. This edge case addresses the question, “What is the worst delay in trial readout one can expect from employing unequal randomization (the worst case being that unequal randomization does not make the trial more attractive, hence enrollment rate stays the same)?”
  • Accelerated accrual: assume unequal randomization will accelerate enrollment to such a rate that the original trial timeline of $\sim$21.7 months is unaffected. This edge case addresses the question, “If the same number of patients is to be enrolled, then how much faster does the enrollment need to be for a trial under unequal randomization to avoid any delay in trial readout?”
  • Increased enrollment: assume unequal randomization will have no impact on enrollment rate (22 patients/month), but additional patients are enrolled at such numbers that the original trial timeline of $\sim$21.7 months is unaffected. This edge case addresses the question, “If unequal randomization does not make the trial more attractive for patient participation (and hence enrollment rate stays the same), then how many additional patients need to be enrolled for a trial under unequal randomization to avoid any delay in trial readout?”
@@
Note that event sizes in Table 1 were determined empirically via simulations. Had Rubinstein’s equation been used instead, similar event sizes would have been obtained (134 for 3:2, 141 for 2:1). However, had Schoenfeld’s equation been used, notably larger event sizes would have been suggested to maintain 80% power (138 for 3:2, 149 for 2:1), which in turn would have exaggerated operating characteristics of the alternative trial designs (i.e., longer trial duration, larger sample size or higher enrollment rates needed to avoid delay).
@@ 4.2 More general cases
We now extend our exercise to cover a wider range of scenarios: hazard ratio (HR) 0.5, 0.6, 0.7, and 0.8; control median (CM) survival 6, 12, and 24 months; and event-patient ratio (EPR) 0.5, 0.6, 0.7, and 0.8. For each “original” trial design employing 1:1 randomization and defined by a unique triplet (HR, CM, EPR), we use the same combination of event size ($d$), sample size ($n$), and accrual rate as was done for the simulations described in Section 2.4. We assume that dropout follows an exponential distribution such that patients have a 1% probability of dropping out every 12 months. Finally, we identify six alternative trial designs by following the approach in Section 4 (“3:2, Prolonged trial”, “3:2, Accelerated accrual”, etc.). Results across all scenarios are summarized in Table 2. (For case-by-case results, see Tables S1–S3.)
Table 2: Comparing number of events and trial duration under 3:2 and 2:1 versus 1:1 randomization in trials with 80% power and various hazard ratios (HR), control medians (CM), and event-patient ratios (EPR). Since under the alternative hypothesis, unequal randomization (without increasing accrual rate or number of patients) leads to longer trials, we also consider the required change in accrual rate (without increasing number of patients) or number of patients (without increasing accrual rate) to avoid any trial delay.
@@
No. of events Minor difference (<$10\%$) Minor difference (<$10\%$), except when treatment effect is weaker and EPR is large (HR $\geq 0.7$, EPR $\geq 0.8$). In this case, no. of events increases >10%.
@@
Trial duration assuming same accrual rate and no. of patients Minor increase (<$10\%$), except when treatment effect is strong, survival is long, and EPR is large (HR $\leq 0.5$, CM $\geq 12$, EPR $\geq 0.8$). In this case, trial duration increases >10%. Minor increase (<$10\%$) when EPR is small (EPR $\leq 0.6$). Increase >$10\%$ otherwise.
@@
— Accrual rate to avoid increase in trial duration Moderate increase (<$25\%$) when survival is short or EPR is small (CM $\leq 6$ or EPR $\leq 0.6$). Increase >$25\%$ otherwise. Increase >$25\%$, except when survival is short and EPR is small (CM $\leq 6$, EPR $\leq 0.6$). In this case, accrual rate increase is moderate (<25%).
@@
— No. of patients to avoid increase in trial duration Minor increase (<$10\%$) Increase >$10\%$, except when treatment effect is strong and EPR is small (HR $\leq 0.6$, EPR $\leq 0.7$). In this case, increase no. of patients is minor (<10%).
@@
Compared to 1:1, 3:2 requires similar number of events to achieve 80% power. In most cases, 3:2 leads to a minor increase in trial duration (assuming the same accrual rate and number of patients), which can be avoided by a moderate increase in accrual rate or minor increase in number of patients. A notable exception is when survival is long and EPR is large (CM $\geq 12$, EPR $\geq 0.7$). In this case, a >25% increase in accrual rate is needed to avoid any trial delay. Trial delay may even be impossible to avoid in some cases; when EPR is 0.8, no amount of increase in accrual rate may offset the slower rate of accumulating events under 3:2 randomization.
@@
2:1 also requires similar number of events as 1:1 in many cases. However, there are some cases when 2:1 requires a >10% increase in number of events (e.g., HR=0.8). 2:1 leads to >10% increase in trial duration when EPR is large (EPR $\geq 0.7$). Trial delay can be difficult to avoid; in many cases, a >25% increase in accrual rate or >10% increase in number of patients is needed. And more often than 3:2, trial delay may be unavoidable. Thus, 2:1 is more likely than 3:2 to result in a notable increase in trial duration, and a greater increase in accrual rate or number of patients is needed to avoid it.
@@
It is worth noting that, in some cases where HR and EPR are small, we do see unequal randomization actually requiring less events than 1:1 randomization. However, the decrease in number of events is always small—2 at most. Therefore, consistent with our observation from Figure 3, 1:1 randomization may not maximize power, but it experiences little power loss compared to 3:2 and 2:1 randomization.
@@
As an important final aside, the fact that Schoenfeld’s approximation is less accurate under unequal randomization may cause us to wonder whether we should stop conducting event-driven survival trials all together; part of the lure of this longstanding practice has been the thought that power depends on the underlying HR, but not some of the other design parameters. We see from Tables S1–S3 that indeed, for trial designs under 3:2 and 2:1 randomization, the required number of events (and therefore power) depends to some extent on the control median survival and EPR. However, we also see that this dependency is relatively weak, much weaker compared to the dependency on the hazard ratio (i.e., changes to $d$ as a function of CM or EPR are much smaller compared to changes to $d$ as a function of HR).
@@
Our recommendation is therefore two-fold. First, randomized survival trials with LRT as the primary analysis should continue to be event-driven. Second, for trials employing unequal randomization, the required number of events should be calculated using a formula-based approach (e.g., Rubinstein or Schoenfeld) as a first iteration only, then subsequently fine-tuned using empirical simulations.11, 32 Alternatively, accurate sample size formulae that rely on numerical integration (and that were originally proposed to accommodate for more complex settings such as non-exponential survival, non-proportional hazards, and treatment switching) can also be considered.12, 33, 34, 35, 36 Our recommendation is consistent with the current practice for trials that anticipate non-proportional hazards (NPH) at the time of trial design. Even though it is well known that under NPH, power depends on many design parameters, these trials remain event-driven. How some have instead adapted is by considering potential NPH when calculating the required number of events (see Checkmate-27437 and Keynote-96638 for example).